OpenStudy (anonymous):

softball: jhon 15 scheduling the games for a softball league to find the number of game he needs to schedule , he uses the equation of 1/2 n squared-1/2 n where g represent the number of games needed for each team to play. Each other team exactly once and n represents the number of teams. write the equation in factor form. how many games are needed in 7 teams to play each other exactly 3 times?

mathslover (mathslover):

\[\Huge{\mathbb{Welcome}\textbf{To }\mathbb{Open}\textbf{Study}}\] I am sure that this will be helpful to you : http://mathnmind.com/PDF%20Files/Algebra1/Textbook/chap09.pdf Just have a look if u r still having problem ? You can ask here further ! thanks !

OpenStudy (anonymous):

\[g = (1/2)n^{2} - (1/2)\]